# Fine-Grained ECG-Text Contrastive Learning via Waveform Understanding Enhancement

### v1

@@ Abstract
Electrocardiograms (ECGs) are essential for diagnosing cardiovascular diseases. While previous ECG-text contrastive learning methods have shown promising results, they often overlook the incompleteness of the reports. Given an ECG, the report is generated by first identifying key waveform features and then inferring the final diagnosis through these features. Despite their importance, these waveform features are often not recorded in the report as intermediate results. Aligning ECGs with such incomplete reports impedes the model’s ability to capture the ECG’s waveform features and limits its understanding of diagnostic reasoning based on those features. To address this, we propose FG-CLEP (Fine-Grained Contrastive Language ECG Pre-training), which aims to recover these waveform features from incomplete reports with the help of large language models (LLMs), under the challenges of hallucinations and the non-bijective relationship between waveform features and diagnoses. Additionally, considering the frequent false negatives due to the prevalence of common diagnoses in ECGs, we introduce a semantic similarity matrix to guide contrastive learning. Furthermore, we adopt a sigmoid-based loss function to accommodate the multi-label nature of ECG-related tasks. Experiments on six datasets demonstrate that FG-CLEP outperforms state-of-the-art methods in both zero-shot prediction and linear probing across these datasets.
Fine-Grained ECG-Text Contrastive Learning via Waveform Understanding Enhancement
Haitao Li1, Che Liu2, Zhengyao Ding1, Ziyi Liu3, Zhengxing Huang1 1Zhejiang University, 2Imperial College London, 3Transtek Medical Electronics Co., Ltd lihaitao@zju.edu.cn, zhengxinghuang@zju.edu.cn
1 Introductin
Electrocardiograms (ECGs) are important, non-invasive tools used in clinical settings to detect heart rhythm disorders Sahoo et al. (2020); Rath et al. (2021); Ayano et al. (2022). ECG self-supervised learning, which uses unlabeled data, has become a promising solution to the limitations of traditional methods that rely on annotated data and require expert knowledge. It can be broadly classified into two types: comparative self-supervision Chen et al. (2020, 2021); Wang et al. (2023); Eldele et al. (2021), which focuses on learning ECG features by distinguishing between augmented positive and negative samples, and generative self-supervision Zhang et al. (2022a); Hu et al. (2023); Na et al. (2024); Zhang et al. (2022b), which aims to reconstruct original ECG signals from masked versions. However, these methods often require annotated samples and struggle with classes not present in the fine-tuning stages.
Figure 1: (a) Doctors typically make diagnoses based on waveform features. (b) Using LLMs to recover waveform features from diagnosis faces two challenges: hallucinations and the non-bijective relationship
@@ 1 Introductin
Significant efforts inspired by vision-text multimodal contrastive learning CLIP Radford et al. (2021) have been made to address such zero-shot challenges. Li et al. (2024); Liu et al. (2024a); Lalam et al. (2023) learn ECG representations by pulling embeddings of paired ECG signals and their corresponding reports while pushing them from unpaired reports, enabling zero-shot prediction. MERL Liu et al. (2024b) further enhances these representations with uni-modal alignment and employs the CKEPE pipeline during inference to generate more descriptive prompts using LLMs. However, enhancing textual prompts only at the inference stage can’t fully leverage the capabilities of LLMs. In contrast, ESI Yu et al. (2024) enhances ECG reports during training by integrating a retrieval-augmented generation (RAG) pipeline that leverages LLMs and external medical knowledge to produce detailed textual descriptions.
However, none of the aforementioned methods noted the missing waveform features in ECG reports. As shown in Figure 1(a), when given an ECG, doctors typically first recognize key waveform features and then make a final diagnosis based on these features. However, these intermediate waveform features are often not recorded in the report, despite being crucial for the model to understand the diagnostic process. Our objective is to recover these omitted features. Notably, simply augmenting the report using LLMs Yu et al. (2024), as shown in Figure 1(b), is unreliable due to: (1) the hallucination of medical LLMs Huang et al. (2023); Günay et al. (2024) and (2) the non-bijective relationship between ECG waveform features and diagnoses Jin (2018), where a single disease may present different waveform characteristics across individuals.
In this study, we propose a novel training strategy FG-CLEP to address the aforementioned challenge. It consists of the following three steps: (1) training the CLEP model using contrastive learning on original ECG-report pairs, (2) generating potential waveform features based on the original report using LLMs and validating them with CLEP, and (3) continuing to train the CLEP model with this augmented report containing waveform features to obtain the final FG-CLEP model. The key is to use the coarse CLEP model to calculate the similarity between the ECG signal and the potential waveform features generated by LLMs. High-probability features are treated as ground truth and incorporated into the reports, which are then used for FG-CLEP training. Thus, this approach not only addresses the non-bijective relationship problem but also provides a mechanism for error correction when LLMs exhibit hallucination issues.
Additionally, ECG data possess unique characteristics compared to images, making direct reuse of the CLIP architecture Radford et al. (2021) suboptimal. First, most ECGs are normal Thai et al. (2017); Yogarajan et al. (2021), leading to similar report semantics across patients. However, previous methods Radford et al. (2021); Liu et al. (2024b, a) treat only ECG-report pairs from the same patient as positive samples, causing frequent false negatives during pre-training. Thus, we introduced a semantic similarity matrix to measure the semantic similarity between different reports, enabling the identification and elimination of false negatives. Second, ECG tasks typically require multi-label classification, but prior approaches use a softmax-based Info-NCE loss Oord et al. (2018), which is unsuitable for such tasks. To address this, we replaced Info-NCE with a sigmoid-based loss to support multi-label scenarios.
Figure 2: Training Porcess of FG-CLEP. (1) training the CLEP model using contrastive learning on original ECG-report pairs, (2) generating potential waveform features based on the original report using LLMs and validating them with CLEP, and (3) continuing to train the CLEP model with this augmented report containing waveform features to obtain the final FG-CLEP model.
@@
We validate our proposed FG-CLEP on six ECG multi-label classification datasets in both zero-shot and linear probing, demonstrating competitive performance compared to state-of-the-art methods in both settings. Overall, our contributions are threefold:
  • We propose a novel training pipeline to capture fine-grained waveform features. Since LLM-generated waveform features are unreliable due to hallucination and the non-bijective relationship, our pipeline validates these features by measuring their similarity to the ECG signal, generating a refined report for continued training.
  • We design a model tailored for ECG multi-label nature and frequent false negatives. We implement a sigmoid-based loss function and introduce a semantic similarity matrix to measure the similarity of reports across patients, guiding contrastive learning and effectively mitigating false negatives.
@@
  • Experimental results indicate that the model, pre-trained on MIMIC-ECG using our FG-CLEP framework, outperforms state-of-the-art methods in both zero-shot prediction and linear probing across six datasets, including PTB-XL, CPSC2018, and CSN.
@@ 2 Related Work
ECG Self-Supervised Learning Self-supervised learning in ECG analysis has primarily been explored through two paradigms: Contrastive self-supervision Chen et al. (2020, 2021); Wang et al. (2023); Eldele et al. (2021), which typically involves augmenting the same ECG signal into two different views as positive samples, while different ECG signals serve as negative samples; Generative self-supervision Zhang et al. (2022b); Hu et al. (2023); Zhang et al. (2022a); Na et al. (2024), which first masks a portion of the ECG signal and then attempts to recover the masked part using the unmasked portion. Unlike previous ECG self-supervised methods that rely on annotations and struggle with unseen classes during fine-tuning, our FG-CLEP enables direct zero-shot prediction on downstream tasks.
@@
ECG-Report Contrastive Learning Recently, inspired by the strong zero-shot ability of image-caption multimodal contrastive learning methods like CLIP Radford et al. (2021), significant efforts have been made in ECG-Report contrastive learningLi et al. (2024); Liu et al. (2024b, a); Yu et al. (2024); Lalam et al. (2023). Similar to CLIP Radford et al. (2021), Li et al. (2024); Liu et al. (2024a); Lalam et al. (2023) learns ECG representations by pulling ECGs with their paired reports while pushing them from unpaired reports. MERL Liu et al. (2024b) further introduces uni-modal alignment and employs the CKEPE pipeline at inference to generate more descriptive prompts via LLMs. However, enhancing textual prompts only during inference creates a distribution mismatch between training and testing text. In contrast, ESI Yu et al. (2024) enhances ECG reports during training using a retrieval-augmented generation (RAG) pipeline, integrating LLMs and external medical knowledge for more detailed descriptions.
@@
Despite these advances, existing methods overlook the absence of fine-grained waveform features in ECG reports. Additionally, due to medical LLM hallucinations Huang et al. (2023); Günay et al. (2024) and the variability of waveform features for the same disease across patients Jin (2018), relying solely on LLMs Yu et al. (2024) for report augmentation is unreliable. To address these challenges, we propose the FG-CLEP training process.
@@
False Negatives in Contrastive Learning Traditional multi-modal contrastive learning Radford et al. (2021) assumes that only images and captions from the same record are positive pairs. However, this assumption often fails in the ECG domain, where most ECGs are normal, and abnormalities typically involve common diseases, leading to frequent false negatives. There have been several attempts to address this issue Lavoie et al. (2024); Jiang et al. (2023b); Sun et al. (2023); Li et al. (2023); Wang et al. (2022). Some approaches Jiang et al. (2023b); Li et al. (2023) attempt to add a regularization term to mitigate false negatives. Others Sun et al. (2023); Wang et al. (2022) introduce a matrix to measure the similarity between different reports, guiding contrastive learning to identify and address false negatives. In this paper, we explore the application of the latter approach in the ECG multi-modal contrastive learning domain.
@@ 3 Method
In this section, we present the technical details of the FG-CLEP framework, which includes a novel pre-training pipeline that recovers missing waveform features in reports and helps the model learn fine-grained ECG representations, as illustrated in Figure 2, as well as a specific model architecture tailored for ECG, shown in Figure 3. This architecture is designed to address the frequent false negatives in ECG-Report pairs and the multi-label nature of ECG downstream tasks. It consists of the following components: (1) ECG and text encoders that extract embeddings, (2) a Semantic Similarity Matrix that measures the similarity of reports from different patients, and (3) $L_{sig}$ and $L_{fnm}$ that train the entire model.
3.1 Model Architecture
As illustrated in Figure 3, FG-CLEP consists of one ECG encoder and one text encoder.
ECG Encoder. We encode ECG signals into embeddings $\mathbf{e}\in\mathbb{R}^{D}$ using an ECG encoder $E_{\text{ecg}}$. A projection head then maps raw embeddings to $\mathbf{e}_{p}\in\mathbb{R}^{P}$.
$\mathbf{e}$ $=E_{\text{ecg}}(x_{\text{ecg}})$ (1a)
$\mathbf{e}_{p}$ $=f_{e}(\mathbf{e})$ (1b)
where $f_{e}$ is the projection head of the ECG encoder.
Text Encoder We create clinically meaningful text embeddings $\mathbf{t}\in\mathbb{R}^{M}$ by a text encoder. We project them to $\mathbf{t}_{p}\in\mathbb{R}^{P}$ as
$\mathbf{t}$ $=E_{\text{txt}}(x_{\text{txt}})$ (2a)
$\mathbf{t}_{p}$ $=f_{t}(\mathbf{t})$ (2b)
where $f_{t}$ is the projection head and $E_{\text{txt}}$ denotes the text encoder. This gives the same embedding dimension $P$ as the ECG encoder, suitable for contrastive learning.
Semantic Similarity Matrix False negatives in the pre-training phase arise from the assumption that ECGs and reports from different patients are unmatched. However, due to the prevalence of common diagnoses, many ECGs exhibit similar symptoms, leading to reports with similar semantics. To address this, we introduce a Semantic Similarity Matrix similar to Sun et al. (2023); Wang et al. (2022) to measure the similarity of reports from different patients during the pre-training phase.
We denote an ECG-report dataset as $D=\{(x_{\text{ecg}_{i}},x_{\text{txt}_{i}})\mid i\in[0,n)\}$, where $(x_{\text{ecg}_{i}},x_{\text{txt}_{i}})$ represents a sample with paired ECG-report content. The ECG and text signals are encoded into $(\mathbf{e}_{p},\mathbf{t}_{p})$ as disscussed above, and the semantic similarity matrix $\mathbf{S}\in\mathbb{R}^{N\times N}$ is defined as follows:
$\mathbf{S}_{ij}=\text{sim}(\mathbf{t}_{pi},\mathbf{t}_{pj})$
where $\text{sim}(\cdot,\cdot)$ is the cosine similarity function.
$\text{sim}(\mathbf{t}_{pi},\mathbf{t}_{pj})=\frac{\mathbf{t}_{pi}\cdot\mathbf{t}_{pj}}{\|\mathbf{t}_{pi}\|\|\mathbf{t}_{pj}\|}$
This results in a similarity matrix $\mathbf{S}$, where each element $\mathbf{S}_{ij}$ indicates the similarity between the $i$-th and $j$-th report embeddings. The Semantic Similarity Matrix guides the contrastive learning process and reduces false negatives by considering the semantic similarities across different pairs. We give an intuitive visualization of the semantic similarity matrix in Section 5.1.
Loss Function The loss function of our FG-CLEP framework consists of two components: $L_{\text{sig}}$ and $L_{\text{fnm}}$. $L_{\text{sig}}$ represents a sigmoid-based loss function tailored to adapt to downstream ECG multi-label classification tasks. $L_{\text{fnm}}$ addresses the issue of false negatives commonly observed in ECG datasets due to the prevalence of common diagnoses.
Sigmoid-based Contrastive Loss $L_{\text{sig}}$ Traditional multimodal contrastive learning often relies on the InfoNCE loss Oord et al. (2018), which involves a softmax operation. However, for ECG downstream tasks, which are mostly multi-label classification problems, using the softmax operation is suboptimcal. To address this problem, we propose using a sigmoid-based loss during the pre-training phase. Specifically, we leverage the sigmoid loss Zhai et al. (2023), which is known for its memory efficiency compared to InfoNCE loss Oord et al. (2018). Here, we use it for its disuse of the softmax operation, making it more suitable for downstream ECG multi-label classification tasks.
The $L_{\text{sig}}$ is defined as:
$L_{\text{sig}}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}\log\left(\frac{1}{1+e^{-z_{ij}(-t\cdot\text{sim}(\mathbf{e}_{pj},\mathbf{t}_{pj})+b)}}\right)$
Where $B$ represents the batch size, $z_{ij}$ denotes the match between a given ECG and report input, equaling 1 if they are paired and -1 otherwise. At initialization, the significant imbalance caused by numerous negatives dominates the loss, resulting in large initial optimization steps aimed at correcting this bias. To mitigate this issue, an additional learnable bias term $b$, similar to the temperature $t$, is used. This ensures that the training starts approximately close to the prior, avoiding the need for substantial over-correction.
Figure 3: The Overview of Model Architecture.
False Negative Mitigation Loss $L_{\text{fnm}}$ Due to the long-tail distribution of ECG labels, most electrocardiograms have similar reports. Simply aligning the embeddings of paired electrocardiograms with their corresponding reports while distinguishing them from unpaired reports results in false negatives. To address this issue, we incorporate the semantic similarity matrix $S$ obtained above into the loss function to guide contrastive learning, similar to Sun et al. (2023). This matrix captures the semantic similarity of reports from different patients, allowing us to identify and correct false negative samples.
The $L_{\text{fnm}}$ is defined as:
$L_{\text{fnm}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}\left|\text{sim}(\mathbf{e}_{pi},\mathbf{t}_{pj})-\mathbf{S}_{ij}\right|$
where $B$ represents the batch size, $\mathbf{S}_{ij}$ is the semantic similarity between the $i$-th and $j$-th report embeddings, and the ||$\cdot$|| ⋅ | represents the L1-distance. We alleviates the false negative problem by reducing the L1 distance between $\text{sim}(\mathbf{e}_{pi},\mathbf{t}_{pj})$ and $\mathbf{S}_{ij}$.
Combined Loss Function The final loss function combines $L_{\text{sig}}$ and $L_{\text{fnm}}$:
$L=L_{\text{sig}}+\lambda L_{\text{fnm}}$
where $\lambda$ is a hyperparameter that controls the trade-off between the two loss components.
In conclusion, by integrating the semantic similarity matrix $\mathbf{S}$ and using a sigmoid-based contrastive loss, our pretraining loss function mitigates the frequent false negatives and better adapts to downstream ECG multi-label tasks when performing zero-shot prediction, leading to improved learning of meaningful representations of ECG signals.
Inference During the inference phase, given an ECG and a text description, we calculate their similarity as described above and use the sigmoid function to determine the probability. The probability score $P$ for the matching between the ECG embedding $\mathbf{e}_{p}$ and the text embedding $\mathbf{t}_{p}$ is computed as follows:
$P=\sigma(\text{sim}(\mathbf{e}_{p},\mathbf{t}_{p}))$
where $\sigma$ denotes the sigmoid function. This probability score indicates how likely it is that the ECG and the text description are related. The higher the score, the more confident we are in the match.
3.2 Training Process

### v2

# Fine-grained Contrastive Learning for ECG-Report Alignment with Waveform Enhancement

@@ Abstract
Electrocardiograms (ECGs) are essential for diagnosing cardiovascular diseases. However, existing ECG-Report contrastive learning methods focus on whole-ECG and report alignment, missing the link between local ECG features and individual report tags. In this paper, we propose FG-CLEP (Fine-Grained Contrastive Language ECG Pre-training), which achieves fine-grained alignment between specific ECG segments and each tag in the report via tag-specific ECG representations. Furthermore, we found that nearly 55% of ECG reports in the MIMIC-ECG training dataset lack detailed waveform features, which hinders fine-grained alignment. To address this, we introduce a coarse-to-fine training process that leverages large language models (LLMs) to recover these missing waveform features and validate the LLM outputs using a coarse model. Additionally, fine-grained alignment at the tag level, rather than at the report level, exacerbates the false negative problem, as different reports may share common tags. To mitigate this, we introduce a semantic similarity matrix to guide the model in identifying and correcting false negatives. Experiments on six datasets demonstrate that FG-CLEP significantly improves fine-grained alignment, outperforming state-of-the-art methods in both zero-shot prediction and linear probing. Meanwhile, the fine-grained reports we generate also enhance the performance of other methods. Our code and data are available at: https://anonymous.4open.science/r/FG-CLEP-3454.
1 Introduction
Electrocardiograms (ECGs) are essential non-invasive tools for detecting cardiac rhythm disorders in clinical practice (Sahoo et al., 2020; Rath et al., 2021; Ayano et al., 2022). Recently, self-supervised learning (SSL) has emerged as a promising paradigm for ECG representation learning, alleviating the reliance on large-scale annotated data and expert knowledge. Existing ECG SSL approaches can be broadly categorized into comparative methods (Chen et al., 2020; 2021; Wang et al., 2023; Eldele et al., 2021) and generative methods (Zhang et al., 2022a; Hu et al., 2023; Na et al., 2024; Zhang et al., 2022b). However, most of these methods still struggle with unseen classes in zero-shot scenarios.
@@ 1 Introduction
Inspired by the success of multimodal contrastive learning such as CLIP (Radford et al., 2021), recent studies have explored ECG–report alignment to enable zero-shot prediction. For instance, Li et al. (2024); Liu et al. (2024a); Lalam et al. (2023) align ECG signals with paired textual reports, while MERL (Liu et al., 2024b) enhances such representations with uni-modal alignment and descriptive prompts from LLMs. ESI (Yu et al., 2024) further integrates retrieval-augmented generation (RAG) (Ni et al., 2025; Gao et al., 2023) pipelines to enrich ECG reports with external medical knowledge. These efforts demonstrate the potential of ECG–text alignment for zero-shot learning.
Despite recent advances, current methods predominantly focus on aligning entire ECGs with their corresponding reports, neglecting the fine-grained relationship between local ECG features and individual report tags. To achieve fine-grained alignment, we identify and address three key challenges.
1.Fine-grained Alignment Architecture: Most existing approaches align entire ECGs with whole reports, but they fail to capture patch-level ECG embeddings and tag-specific report embeddings, both of which are critical for fine-grained alignment.
2.Missing Waveform Features in Reports: We observe that nearly 55% of ECG reports in the MIMIC-ECG dataset—one of the largest ECG-Report datasets—lack detailed waveform features, which hampers fine-grained alignment. In clinical practice, physicians often begin by identifying key waveform patterns in an ECG before formulating a diagnosis. However, many physicians do not explicitly record these features in the reports, resulting in a significant portion of reports lacking important waveform information. Recovering these missing features using large language models (LLMs) is challenging for two main reasons: (1) the hallucination problem inherent in LLMs (Huang et al., 2023; Günay et al., 2024; Zhang et al., 2025), and (2) the non-bijective relationship between waveform features and diagnostic outcomes, where the same disease may manifest with different waveform patterns (Jin, 2018). As a result, relying solely on LLMs to augment reports, as attempted by Yu et al. (2024), proves to be unreliable.
3.False Negative Challenge: Fine-grained alignment at the tag level, rather than at the report level, exacerbates the false negative problem. This is because different reports may share common tags, leading to potential misalignments.
In this study, we propose FG-CLEP to address the aforementioned challenges. Rather than aligning the entire ECG embedding with the report embedding, we perform alignment at the ECG patch level and the individual tag level in the report. Specifically, we obtained tag-specific ECG representations, where different tags in the report serve as queries to attend to ECG patches (treated as keys and values) through cross-attention. To address missing waveform features in reports and generate fine-grained reports, we propose a coarse-to-fine training process. First, we train a coarse CLEP model using contrastive learning on the original ECG-report pairs. Then, we use LLMs to generate potential waveform features from the reports, which are validated using CLEP. Finally, we integrate these validated features into the reports and continue training the CLEP model to obtain the final FG-CLEP model. This approach resolves the non-bijective relationship between waveform features and diagnoses and corrects errors from LLM hallucinations. Lastly, to address the false negative problem exacerbated by fine-grained tag-level alignment, we introduce a semantic similarity matrix. This matrix computes the similarity between tags and is used during contrastive learning to guide the model in identifying and correcting false negatives.
@@
We validate our proposed FG-CLEP on six ECG multi-label classification datasets in both zero-shot and linear probing, the results demonstrate that FG-CLEP significantly improves fine-grained alignment, outperforming state-of-the-art methods in both zero-shot prediction and linear probing. Meanwhile, the fine-grained reports we generate also enhance the performance of other methods. Overall, our contributions are threefold:
  • We propose a new alignment architecture, enabling fine-grained alignment between ECG segments and report tags through tag-specific ECG representations, capturing detailed ECG-report relationships.
  • We introduce a coarse-to-fine training process using LLMs to recover missing waveform features and validate them with a coarse CLEP model, addressing non-bijective relationships and LLM hallucinations.
  • We present a semantic similarity matrix to mitigate false negatives in ECG-tag pairs, guiding contrastive learning to correct misalignments.
@@
  • Experimental results show that FG-CLEP, pre-trained on MIMIC-ECG, outperforms state-of-the-art methods in zero-shot prediction and linear probing across six datasets, including PTB-XL, CPSC2018, and CSN.
@@ 2 Related Work
ECG-Report Contrastive Learning Recently, inspired by the strong zero-shot ability of image-caption multimodal contrastive learning methods like CLIP (Radford et al., 2021), significant efforts have been made in ECG-Report contrastive learning (Li et al., 2024; Liu et al., 2024b; a; Yu et al., 2024; Lalam et al., 2023). Similar to CLIP (Radford et al., 2021), Li et al. (2024); Liu et al. (2024a); Lalam et al. (2023) learns ECG representations by pulling ECGs with their paired reports while pushing them from unpaired reports. MERL (Liu et al., 2024b) further introduces uni-modal alignment and employs the CKEPE pipeline at inference to generate more descriptive prompts via LLMs. However, enhancing textual prompts only during inference creates a distribution mismatch between training and testing text. In contrast, ESI (Yu et al., 2024) enhances ECG reports during training using a retrieval-augmented generation (RAG) pipeline, integrating LLMs and external medical knowledge for more detailed descriptions.
@@
Despite these advances, existing methods all perform contrastive learning at the whole-ECG and report level, and overlook the absence of fine-grained waveform features in ECG reports. To address these challenges, we propose the FG-CLEP.
Fine-Grained Contrastive Learning Recent studies in medical imaging have demonstrated the benefits of fine-grained alignment. Methods such as MedFILIP (Liang et al., 2025), fVLM (Shui et al., 2025) incorporate entity-level or region-level supervision by leveraging structured information extracted from reports or by segmenting images into anatomical regions. These approaches enable more precise matching between local visual features and textual descriptions, resulting in improved downstream performance, especially in zero- and few-shot settings.
However, the application of fine-grained contrastive learning in the ECG domain remains largely unexplored. To bridge this gap, we propose to model fine-grained correspondences between characteristic ECG waveform segments and specific textual patterns in the reports, aiming to capture subtle diagnostic cues that may be overlooked in global-level alignment.
@@
False Negatives in Contrastive Learning Traditional multi-modal contrastive learning (Radford et al., 2021) assumes that only images and captions from the same record are positive pairs. However, this assumption often fails in the ECG domain, where most ECGs are normal, and abnormalities typically involve common diseases, leading to frequent false negatives. Furthermore, fine-grained alignment at the tag level exacerbates this issue, as different reports may share common tags. There have been several attempts to address this issue (Lavoie et al., 2024; Jiang et al., 2023b; Sun et al., 2023; Li et al., 2023; Kim et al., 2025). Some approaches (Jiang et al., 2023b; Li et al., 2023) attempt to add a regularization term to mitigate false negatives. Others (Sun et al., 2023; Wang et al., 2022) introduce a matrix to measure the similarity between different reports, guiding contrastive learning to identify and address false negatives. In this paper, fine-grained alignment introduces a more pronounced false negative problem, as different reports may share common tags.
Figure 1: Fine-grained Alignment Architecture with Tag-Specific ECG Representation.
@@ 3 Method
3.1 Fine-Grained Alignment
In this section, we describe how we achieve fine-grained alignment between specific ECG segments and tags in clinical reports. Given an ECG-report dataset $\mathcal{D}=\{(\mathbf{X}_{i},\mathcal{R}_{i})\}_{i=1}^{M}$, each report $\mathcal{R}_{i}$ consists of multiple clinically meaningful tags, such as arrhythmia, myocardial infarction, atrial fibrillation, etc. For each tag, our objective is to identify and align the corresponding ECG segment(s) that reflect this clinical finding. To achieve this, we first perform ECG patch-level encoding and report tag-level encoding to obtain fine-grained representations of both the ECG signals and the report tags.
ECG Patch-Level Encoding.
To obtain fine-grained ECG representations, we employ ViT to encode patch-level features from the ECG signal. Given a 12-lead ECG $\mathbf{X}_{i}\in\mathbb{R}^{12\times T}$, where $L=12$ is the number of leads and $T$ is the signal length, we independently divide each lead into $N_{\text{lead}}$ non-overlapping segments (patches) along the temporal axis, with each patch of length $\Delta T=T/N_{\text{lead}}$. This yields a total of $N=12\times N_{\text{lead}}$ patches:
$\mathbf{X}_{i}\rightarrow\left\{\mathbf{x}_{i}^{p}\right\}_{p=1}^{N},\quad\mathbf{x}_{i}^{p}\in\mathbb{R}^{1\times\Delta T}.$ (1)
All $N$ patches are passed through the ECG encoder resulting in a sequence of patch-level embeddings:
$\mathbf{H}_{i}=\left[\mathbf{h}_{i}^{1},\ldots,\mathbf{h}_{i}^{N}\right]=\mathrm{ECGEncoder}\left(\left[\mathbf{x}_{i}^{1},\ldots,\mathbf{x}_{i}^{N}\right]\right)\in\mathbb{R}^{N\times d}.$ (2)
Report Tag-Level Encoding.
Since ECG reports are highly structured, we can simply use commas to split the tags. We denote the set of tags in the $i$-th report as $\mathcal{T}_{i}=\{t_{i}^{j}\}_{j=1}^{m_{i}}$, where $t_{i}^{j}$ represents the $j$-th tag in report $i$, and $m_{i}$ is the number of tags in $\mathcal{R}_{i}$. Each tag $t_{i}^{j}$ is then independently encoded via a text encoder to obtain its embedding representation. This yields the tag-level representation matrix for report $i$:
$\mathbf{Z}_{i}=[\mathbf{z}_{i}^{1},\dots,\mathbf{z}_{i}^{m_{i}}]\in\mathbb{R}^{m_{i}\times d},$ (3)
$\mathbf{z}_{i}^{j}=\mathrm{TextEncoder}(t_{i}^{j}),\quad j=1,\dots,m_{i}.$ (4)
Notably, both the ECG patch embeddings and tag embeddings are projected into the same latent space with dimension $d$. For clarity and conciseness, we omit explicit projection formulas in the equations above.
Tag-Specific ECG Representation.
With the above fine-grained ECG patch embeddings and report tag embeddings, we can now align them at a fine-grained level. Ideally, if explicit annotations mapping each tag to its corresponding ECG patches were available, direct supervised alignment could be applied. However, such fine-grained annotations are typically unavailable in practice. To address this, we propose an automatic alignment mechanism based on cross-attention.
Specifically, for each tag, we use its embedding as the query and the ECG patch embeddings as keys and values. Through the cross-attention mechanism, the model adaptively computes a tag-specific ECG representation by attending to ECG patches according to their relevance to the tag. This enables each tag to aggregate information from the most relevant ECG segments, capturing the fine-grained relationship without requiring extra supervision.
Formally, given a tag embedding $\mathbf{z}_{i}^{j}$ and ECG patch embeddings $\mathbf{H}_{i}$, we compute the tag-specific ECG representation as:
$\mathbf{e}_{i}^{j}=\mathrm{CrossAttn}(\mathbf{z}_{i}^{j},\mathbf{H_{i}})=\sum_{p=1}^{N}\alpha_{j,p}\cdot\mathbf{h}_{i}^{p},$ (5)
$\alpha_{j,p}=\frac{\exp(\langle\mathbf{z}_{i}^{j},\mathbf{h}_{i}^{p}\rangle)}{\sum_{p^{\prime}=1}^{N}\exp(\langle\mathbf{z}_{i}^{j},\mathbf{h}_{i}^{p^{\prime}}\rangle)}.$ (6)
Fine-Grained Contrastive Learning Objective.
For each random batch with size $B$, we sample one tag $z_{i}$ from each report’s tag set $\mathbf{Z}_{i}$, and compute its corresponding tag-specific ECG representation $\mathbf{e}_{i}$. The learning objective is to maximize the similarity between each tag embedding and its corresponding ECG representation, while minimizing the similarity with unpaired ones, similar to SigLIP (Zhai et al., 2023) for its efficiency compared to the original CLIP.
$L_{\text{con}}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}\log\left(\frac{1}{1+\exp\left(-y_{ij}\cdot t\cdot\text{sim}(\mathbf{e}_{i},\mathbf{z}_{j})\right)}\right).$ (7)
where $y_{ij}$ denotes the match between a given ECG and report input(1 if i==j, otherwise -1), $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity, and $t$ is a temperature hyperparameter.
Figure 2: Coarse-to-Fine Training Process of FG-CLEP. The reports with recovered waveform features are referred to as fine-grained reports. Models trained with fine-grained reports are prefixed with FG-
3.2 Coarse-to-Fine Training Process